# Automatic Forward Model Parameterization with Bayesian Inference of Conformational Populations

### v1

@@ III Results/Discussion
In Figure 4, we show the free energy landscape, which is also equivalent to the BICePs score landscape $f_{\xi=0\rightarrow 1}$. The Karplus curve for ${{}^{3}\!J}_{H^{N}C^{\prime}}$ was found to overlap strongly with the results obtained by SVD when using $\phi$ angles from the X-ray crystal structure (Figure S9). Red data points are shown using the experimental $J$-couplings with $\phi$ angles derived from X-ray crystal pose 1UBQVijay-Kumar, Bugg, and Cook (1987).
@@
The joint BICePs score landscape for the six sets of parameters is too complex to visualize. In an attempt to do our best, we constructed a smooth 2-D landscape for each pair of parameters within a set of scalar couplings by training a Gaussian process on the BICePs energy trace using a radial basis function (RBF) kernel. The landscape matches the computed BICePs scores, and shows minima in the correct locations. All BICePs score landscapes for each of the six sets of Karplus coefficients are illustrated in Figure S14.
To demonstrate the transferability across different generative models and validate our parameters, we evaluated the accuracy of the back-calculated scalar couplings using the different sets of Karplus coefficients. In Figure 5, we illustrate how the various sets of parameters derived from different techniques and different structural ensembles exhibit similar performance metrics. Interestingly, applying BICePs-refined Karplus parameters to an ensemble generated by a molecular dynamics simulation (CHARMM22*),Piana, Lindorff-Larsen, and Shaw (2013) some parameter sets are revealed to be more transferable than others. The mean absolute error (MAE) and coefficient of determination ($R^{2}$) for all six types of scalar couplings across different structural ensembles are shown in Figures S15-S17. On average, the BICePs-refined parameters derived from the 2NR2 ensemble (BICePs(2NR2)) give the lowest MAE between experiment and predictions for the CHARMM22* simulated ensemble, closely followed by BICePs(RF2) parameters, whereas Habeck 2005 has the highest due to known difficulties with ${{}^{3}\!J}_{C^{\prime}C^{\beta}}$.
@@
To objectively quantify which parameters produce the best predictions for ubiquitin, we compute BICePs scores, $f_{\xi=0\rightarrow 1}$ for each of the structural ensembles. This score directly relates to the quality of FM parameters and their predictive accuracy at reproducing experimental scalar couplings, while taking into consideration all sources of potential error. Lower BICePs scores indicate better agreement with experiment. Each row in Table 2 corresponds to BICePs scores using all six sets of Karplus coefficients used on different structural ensembles. The lowest score is shown in bold.
BICePs scores, $f_{\xi=0\rightarrow 1}$ were computed to objectively rank the quality of FM parameters and their predictive accuracy at reproducing experimental scalar couplings (Table 2). The leftmost column in Table 2 corresponds to the parameters, where BICePs(1D3Z) are the parameters in Table 1 (set 1), which used 1D3Z ensemble to obtain Karplus coefficients. BICePs score columns, e.g., $f_{\xi=0\rightarrow 1}^{\text{1d3z}}$ corresponds to BICePs scores evaluated for the 1D3Z ensemble. That is, the superscript corresponds to the structural ensemble used as a validation step. BICePs scores, $f$ for each structural ensemble over all sets of parameters, averaged over five independent rounds of validation each. BICePs calculations burned for 1k steps, followed by 50k steps of MCMC sampling.
Note that the BICePs score is an extensive quantity that grows linearly with the number of replicas. For this reason, our results report the reduced BICePs score, $f(\theta)/N_{r}$. We can confirm that the BICePs score, $f_{\xi=0\rightarrow 1}^{\text{1d3z}}=38.14\pm 0.08$ (Table 2) is equivalent (within error) with the most probable landscape basin $f=38.15\pm 0.19$ from sampling the energy landscape, computed as an averarge across four chains; an example for one chain is shown in Figure 4. This is additional evidence of the algorithm’s reliability and quality of the BICePs score, corroborating that the results from variational minimization of the BICePs score and full joint posterior sampling are equivalent.
@@
For both the 2NR2 and CHARMM22* structural ensembles, Bax 1997, BICePs(RF2) and BICePs(1D3Z) parameters give very similar BICePs scores, which suggests robust accuracy of FM parameters in reproducing experimental scalar couplings and the transferability of FM parameters across different prior structural ensembles. Furthermore, when it comes to the CHARMM22* simulated ensemble, the BICePs(2NR2) parameters give the lowest BICePs score. However, it is important to note that the structural ensemble 2NR2 was generated using CHARMM22 force field with additional experimental restraints during simulation.
@@
Ensembles from generative models like RosettaFold2 can be used for parameter refinement.
The booming field of machine learning and artificial intelligence is swiftly transforming the field of modeling structure and dynamics in biological systems. Recent advancements in generative models, such as AlphaFoldJumper et al. (2021), RosettaFoldBaek et al. (2023) and others, have heralded a new era in the accurate prediction of structural ensembles. Leveraging the predictive power of these models as structural priors is expected to help refine ensemble predictions when integrated with similar algorithms to BICePsBrotzakis, Zhang, and Vendruscolo (2023). Here, we have demonstrated that structural ensembles generated from RosettaFold2 (RF2) can be reweighted to better align with experimental measurements, while simultaneously refining Karplus parameters. Validation of these parameters by the BICePs score and other statistics demonstrates improved accuracy across a varity of structural ensembles of ubiquitin.
@@
In the context of model validation, the BICePs score emerges as a superior metric over the traditional $\chi^{2}$ test. Unlike $\chi^{2}$, which presupposes a fixed and known error, BICePs dynamically ascertains the level of uncertainty, thereby providing a more nuanced and accurate measure of model quality.
@@
The Karplus equation, a cornerstone for interpreting NMR spectroscopy data, comes in multiple forms to accommodate the diverse characteristics of molecular structures, from rigid to flexibleMinch (1994). The BICePs algorithm can determine coefficients and their uncertainties for any functional form, including those with additional parameters. Although we do pursue this aim in our current work, it is straightforward to use Bayesian model selection to objectively rank empirical models based on their BICePs scores, while automatically accounting for model complexity, thus providing a balance of model accuracy and parsimony.
Adaptive variance as simulated annealing.
As an alternative approach to determine optimal FM parameters, we propose that future work might utilize an annealing approach, in which the variance parameter $\sigma^{2}$ is akin to the temperature. Initially, a high $\sigma^{2}$ would enable extensive exploration of the parameter space to circumvent local optima. This exploration phase mimics the high-temperature regime in annealing, allowing for a broad search. Subsequently, we suggest a schedule of stepwise reduction in $\sigma^{2}$, similar to cooling in simulated annealing, to gradually narrow the search area and determine the optimum solution. This method balances between wide-ranging search and focused refinement, potentially enhancing search efficiency and robustness in FM parameter optimization.
IV Conclusion
@@
In the quest for accurate forward model predictions, specifically for $J$-coupling, researchers often navigate the vast literature seeking Karplus parameters that align with their specific systems, occasionally settling for less-than-ideal solutions. Our work demonstrates BICePs as a robust tool for determining forward model (FM) parameters by sampling over their full posterior distribution. We used a toy model to demonstrate that variational minimization of the BICePs score is also a valid approach for FM parameter refinement.
We have shown how the BICePs score–the free energy of “turning on” the restraints tethering the forward model predictions to the experimental values–serves as an effective validation metric for FM parameters. Using structural ensembles and experimental data for ubiquitin, BICePs determined six different sets of Karplus coefficients using different types of $J$-coupling measurements, while effectively addressing both random and systematic errors. From these results, one can see how this algorithm can be applied more generally to find other optimal forward model parameters. These advances not only contribute to the refinement of molecular simulations but also hold promise for a wide range of applications within the scientific community, particularly among those analyzing structural dynamics and performing model validation.
@@ Acknowledgements
RMR, TM and VAV are supported by National Institutes of Health grant R01GM123296. This research includes calculations carried out on HPC resources supported in part by the National Science Foundation through major research instrumentation grant number 1625061 and by the US Army Research Laboratory under contract number W911NF-16-2-0189.
Appendix A Variational minimization of the BICePs score to find optimal parameters.
It is likely that the second derivative w.r.t FM parameters would be useful for second-order optimization methods and uncertainty quantification. For the sake of simplicity, we refrain from writing out the second derivative for the Good-Bad model here. However, when $\varphi=1$ the second partial derivatives of the BICePs energy with respect to parameters $\theta_{i}$ and $\theta_{k}$ are:
$\frac{\partial^{2}u}{\partial\theta_{i}\partial\theta_{k}}=N\left[\sum_{j=1}^{N_{d}}-\frac{\partial^{2}g_{j}(X,\theta)}{\partial\theta_{i}\partial\theta_{k}}\frac{(d_{j}-g_{j}(\mathbf{X},\theta))}{\sigma_{j}^{2}}\right.$ (21)
$\left.+\frac{\partial g_{j}(X,\theta)}{\partial\theta_{i}}\cdot\frac{\partial g_{j}(X,\theta)}{\partial\theta_{k}}\frac{1}{{\sigma_{j}}^{2}}\right].$
When $i$ = $k$, the second partial derivative of the energy is just:
$\frac{\partial^{2}u}{\partial\theta^{2}}=N\left[\sum_{j=1}^{N_{d}}\left(-\frac{\partial^{2}g_{j}(X,\theta)}{\partial\theta^{2}}\frac{(d_{j}-g_{j}(\mathbf{X},\theta))}{\sigma_{j}^{2}}\right.\right.$ (22)
$\left.\left.+\left(\frac{\partial g_{j}(X,\theta)}{\partial\theta}\right)^{2}\frac{1}{{\sigma_{j}}^{2}}\right)\right].$
@@
The second partial derivatives of the BICePs score with respect to parameters $\theta_{i}$ and $\theta_{k}$ are:
@@
$\frac{\partial^{2}f(\theta)}{\partial\theta_{i}\partial\theta_{k}}=\bigg{\langle}\frac{\partial^{2}u}{\partial\theta_{i}\partial\theta_{k}}\bigg{\rangle}-\biggl{(}$ $\bigg{\langle}\frac{\partial u}{\partial\theta_{i}}\cdot\frac{\partial u}{\partial\theta_{k}}\bigg{\rangle}-\bigg{\langle}\frac{\partial u}{\partial\theta_{i}}\cdot\frac{\partial u}{\partial\theta_{k}}\bigg{\rangle}^{\prime}\biggr{)},$ (23)
@@
where the notation $\langle\cdot\rangle^{\prime}$ denotes the ensemble average with respect to $\left(\frac{1}{Z}\exp(-u)\right)^{2}$. The first term on the right is the ensemble-averaged second derivative of the energy function $u$ with respect parameters $\theta_{i}$ and $\theta_{k}$ is equation 22 (when $\varphi=1$).
@@
When $i$ = $k$, the second partial derivative of the BICePs score reduces to the difference between the ensemble-averaged second derivative of the energy $u$ and the variance of its first partial derivative:
@@
$\begin{split}\frac{\partial^{2}f}{\partial\theta^{2}}&=\bigg{\langle}\frac{\partial^{2}u}{\partial\theta^{2}}\bigg{\rangle}-\left(\bigg{\langle}\left(\frac{\partial u}{\partial\theta}\right)^{2}\bigg{\rangle}-\bigg{\langle}\frac{\partial u}{\partial\theta}\bigg{\rangle}^{2}\right)\end{split}$ (24)
@@
The calculation is performed using the MBAR free energy estimator for the BICePs score and it’s derivatives, by sampling at several intermediates $\xi=0\rightarrow 1$, which enables accurate estimates of all quantities.
In the case of convex landscapes, one can compute reported uncertainties in the set of optimized FM parameters $\theta$ by estimation of the covariance through inversion of the Hessian (the matrix of second partial derivatives of the BICePs score). Conceptually, this means that the estimated uncertainties in the best-fit values represent the widths of the basins on the BICePs score landscape. For non-convex problems, this can sometimes lead to an under-estimation of uncertainty depending upon the curvature of the basins of the local minima in the BICePs score landscape.
Appendix B Supplemental Information
B.1 Methods
Structural ensembles of ubiquitin

### v2

@@ III Results
In Figure 4, we show the free energy landscape, which is also equivalent to the BICePs score landscape $f_{\xi=0\rightarrow 1}$. The Karplus relation for ${{}^{3}\!J}_{H^{N}C^{\prime}}$ was found to overlap strongly with the results obtained by SVD when using $\phi$ angles from the X-ray crystal structure (Figure S7). Red data points are shown using the experimental $J$-couplings with $\phi$ angles derived from X-ray crystal pose 1UBQVijay-Kumar, Bugg, and Cook (1987).
@@
The joint BICePs score landscape for the six sets of parameters is too complex to visualize. Instead,we constructed a smooth 2-D landscape for each pair of parameters within a set of scalar couplings by training a Gaussian process on the BICePs energy trace using a radial basis function (RBF) kernel with an additive white noise to account for observational uncertainty. The characteristic length scale was bounded within $[0.1,10.0]$ and the signal variance was set to 1.0 and the noise variance was set to $10^{-5}$. The landscape matches the computed BICePs scores, and shows minima in the correct locations. All BICePs score landscapes for each of the six sets of Karplus coefficients are illustrated in Figure S12.
To demonstrate the transferability across different generative models and validate our parameters, we evaluated the accuracy of the back-calculated scalar couplings using the different sets of Karplus coefficients. In Figure 5, we illustrate how the various sets of parameters derived from different techniques and different structural ensembles exhibit similar performance metrics. Interestingly, applying BICePs-refined Karplus parameters to an ensemble generated by a molecular dynamics simulation (CHARMM22*),Piana, Lindorff-Larsen, and Shaw (2013) some parameter sets are revealed to be more transferable than others. The mean absolute error (MAE) and coefficient of determination ($R^{2}$) for all six types of scalar couplings across different structural ensembles are shown in Figures S13-S15. On average, the BICePs-refined parameters derived from the 2NR2 ensemble (BICePs(2NR2)) give the lowest MAE between experiment and predictions for the CHARMM22* simulated ensemble, closely followed by BICePs(RF2) parameters, whereas Habeck 2005 has the highest due to known difficulties with ${{}^{3}\!J}_{C^{\prime}C^{\beta}}$. A visualization of the BICePs-reweighted 2NR2 structural ensemble using BICePs(2NR2) parameters is shown in Figure 6 with higher transparency representing lower populations. Residues are colored by the magnitude of the replica-averaged data deviation from the experiment in ${{}^{3}\!J}_{C^{\prime}C^{\prime}}$, where the color is normalized from blue (lowest deviation) to red (highest deviation).
@@
To quantify which parameters produce the best predictions for ubiquitin, we compute BICePs scores, $f_{\xi=0\rightarrow 1}$ for each of the structural ensembles. This score directly relates to the quality of FM parameters and their predictive accuracy at reproducing experimental scalar couplings, integrating over all sources of error. Lower BICePs scores indicate better agreement with experiment. BICePs scores were computed using eleven intermediate $\xi$-values between 0 and 1, which were optimized using pylambdaopt.Novack et al. (2025) The results are shown in Table 2. Each row in Table 2 corresponds to BICePs scores using all six sets of Karplus coefficients used on different structural ensembles. The lowest score is shown in bold. (Note that the BICePs score is an extensive quantity that grows linearly with the number of replicas. For this reason, our results report the reduced BICePs score, $f(\theta)/N_{r}$. )
The leftmost column in Table 2 corresponds to the parameters, where BICePs(1D3Z) are the parameters in Table 1 (set 1), which used 1D3Z ensemble to obtain Karplus coefficients. BICePs score columns, e.g., $f_{\xi=0\rightarrow 1}^{\text{1d3z}}$ corresponds to BICePs scores evaluated for the 1D3Z ensemble. That is, the superscript corresponds to the structural ensemble used as a validation step. The reported BICePs scores, were averaged over five independent rounds of validation each. BICePs calculations burned for 1k MCMC steps, followed by 50k steps of production-run MCMC sampling.
The BICePs score, $f_{\xi=0\rightarrow 1}^{\text{1d3z}}=38.14\pm 0.08$ (Table 2) is equivalent (within error) with the most probable landscape basin $f=38.15\pm 0.19$ from sampling the energy landscape, computed as an averarge across four chains; an example for one chain is shown in Figure 4. This is additional evidence of the algorithm’s reliability ,and corroborates that the results from variational minimization of the BICePs score and full joint posterior sampling are equivalent.
@@
For both the 2NR2 and CHARMM22* structural ensembles, Bax 1997, BICePs(RF2) and BICePs(1D3Z) parameters give very similar BICePs scores, which suggests robust accuracy of FM parameters in reproducing experimental scalar couplings and the transferability of FM parameters across different prior structural ensembles. For the CHARMM22* simulated ensemble, the BICePs(2NR2) parameters give the lowest BICePs score. A partial explanation for this is that the structural ensemble 2NR2 deposited in the Protein Data Bank was generated using the CHARMM22 force field with additional experimental restraints during simulation.Richter et al. (2007)
Figure 6: A visualization of the 2NR2 conformational ensemble after joint conformational populations and Karplus parameter optimization using BICePs. Structures (144 in total) with higher transparency have lower populations. Residues are colored by the magnitude of the replica-averaged data deviation from the experiment in ${{}^{3}\!J}_{C^{\prime}C^{\prime}}$, where the color is normalized from blue (lowest deviation) to red (highest deviation).
@@
IV Discussion
Ensembles from deep generative models like RosettaFold2 can be used for parameter refinement.
The booming field of machine learning and artificial intelligence is transforming the field of biomolecular modeling at a swift pace. Recent advancements in generative models, such as AlphaFoldJumper et al. (2021), RosettaFoldBaek et al. (2023) and others, have heralded a new era in the accurate prediction of structural ensembles. Leveraging the predictive power of these models as structural priors is expected to help refine ensemble predictions when integrated with similar algorithms to BICePsBrotzakis, Zhang, and Vendruscolo (2023). Here, we have demonstrated that structural ensembles generated from RosettaFold2 (RF2) can be reweighted to better align with experimental measurements, while simultaneously refining Karplus parameters. Validation of these parameters by the BICePs score and other statistics demonstrates improved accuracy across a varity of structural ensembles of ubiquitin.
@@ IV Discussion
In previous work, we have shown how the BICePs score is a better metric model validation than the traditional $\chi^{2}$ test.Raddi et al. (2025) Unlike $\chi^{2}$, which presupposes a fixed and known error, BICePs dynamically learns the complete posterior distribution of all uncertainty parameters, providing a measure of model quality that integrates over all sources of error. Our work here supports this premise.
@@
Although here we optimize Karplus relations with a standard functional form, many possible relations can be used to accommodate the diverse characteristics of molecular structures, from rigid to flexible. Minch (1994) The BICePs algorithm can determine coefficients and their uncertainties for any functional form, including those with additional parameters. Although we do pursue this aim in our current work, it is straightforward to use Bayesian model selection to objectively rank empirical models based on their BICePs scores, while automatically accounting for model complexity, thus providing a balance of model accuracy and parsimony.
Application to Non-Natural and Cyclic Peptides for fine-tuning folding landscapes.
Our framework for forward model optimization has recently been successfully applied to a challenging test case involving a diverse set of non-natural and cyclic $\beta$-hairpin peptidesNguyen, Raddi, and Voelz (2025). In that study, we used BICePs’ forward model optimization approach to refine Karplus parameters for ${{}^{3}\!J}_{H^{N}H^{\alpha}}$ predictions directly against experimental NMR data. With an improved Karplus relation, BICePs reshaped the folding landscapes to better agree with NMR data, enabling more accurate predictions of folding stability for each peptide. This work highlights the resolution enhancement gained by the improved forward model predictions and population reweighting using BICePs, enabling the detection of subtle changes in folding stability—such as those introduced by a sidechain hydrogen- or halogen-bonding group, which alters stability by no more than 2 kJ/mol—offering a reliable pathway for designing foldable non-natural and cyclic peptides.
The BICePs score can be used as a loss function to train neural networks and other differentiable forward models.
A central challenge in physical sciences is integrating domain knowledge with data-driven models. Traditional machine learning often lacks physical consistency or interpretability, while purely physics-based models may miss subtle experimental trends. One way to bridge this gap is to use the BICePs score—a free energy-based Bayesian loss—to train machine learning forward models using ensemble-averaged experimental data. This allows one to parameterize neural networks or other differentiable models in a physically grounded, uncertainty-aware fashion. In this way, BICePs provides a general-purpose learning framework for constructing interpretable, probabilistically-justified forward models for complex observables. This capability may be especially valuable in contexts where observables are too complex to be modeled analytically—such as in chemical shift prediction or other nonlinear structure–observable relationshipsHan et al. (2024); Fang et al. (2021); Sanchez-Garcia et al. (2021).
To demonstrate the viability of using BICePs for neural network training, we constructed a toy model for $J$-coupling prediction. In this model, the backbone dihedral angle $\phi$ was the sole input variable. To appropriately encode the periodicity of angular data, each $\phi$-angle was embedded into two dimensions via a sine–cosine transformation:
$x=\begin{bmatrix}\cos(\phi)\\ \sin(\phi)\end{bmatrix}\in\mathbb{R}^{2}$ (22)
This representation ensures continuity across the periodic boundary at ±$\pm\pi$ and provides a smooth input feature space for learning.
A fully connected feedforward neural network was constructed using this 2D input, consisting of one hidden layer with 200 neurons and Gaussian Error Linear Unit (GELU) activation functions. The final output was a scalar prediction corresponding to the $J$-coupling value. To ensure stable training dynamics compatible with GELU, the weights were initialized using LeCun normal initializationLeCun et al. (2002).
The network was trained using ADAM optimization for 2000 epochs with a learning rate of $10^{-3}$, using the Good-Bad likelihood model with the BICePs score as the loss function. The toy dataset consisted of 5 conformational states, each with 60 synthetic $J$-coupling observables using the Karplus relation $J(\phi)=A\cos^{2}(\phi)+B\cos(\phi)+C$, with the ground truth coefficients $A=6.51$, $B=-1.76$, and $C=1.60$. Our training dataset consisted of 300 $\phi$-angles as NN inputs.
First, we test the performance without adding random or systematic noise to the synthetic experimental J-coupling observables. Figure S16a shows the convergence behavior of the BICePs score across five independent training runs, where a plateau is reached around step 1500. Figure S16b shows the Karplus curves predicted by the five trained networks, along with their mean and standard deviation (blue curve). For comparison, we also include the result of singular value decomposition (SVD) applied directly to the Karplus form under the assumption that backbone $\phi$ angles are known precisely, as would be the case with a high-resolution crystal structure. The parameters from SVD were determined to be $A=6.47\pm 0.002$, $B=-1.75\pm 0.001$, and $C=1.62\pm 0.001$.
Next, we evaluate the performance when adding random or systematic noise to the synthetic experimental $J$-coupling observables ($\sigma_{\text{data}}=0.79$ Hz). Figure S17a–b shows the convergence behavior of the BICePs score and gradient norm across five independent training runs, where convergence begins around 1000 epochs. The BICePs score was determined to be $f_{\xi=0\rightarrow 1}=62.29$ nats. Figure S17c shows the Karplus curves predicted by the five trained networks, along with their mean and standard deviation (blue curve). Comparison against SVD illustrates that BICePs, when used to train neural networks from conformational ensembles alone, recovers predictive performance comparable to analytical fitting based on ground-truth dihedral angles, even in the presence of random and systematic errors.
Convergence of the BICePs score is typically accompanied by a reduction in the posterior uncertainty parameter $\sigma$, which sharpens the likelihood landscape and amplifies gradients with respect to model parameters. Specifically, the derivative of the BICePs score with respect to a parameter $\theta$ contains a term proportional to $1/\sigma^{2}$ (see equations 12 and 13), meaning that as $\sigma\to 0$ the sensitivity of the loss to prediction errors increases.
During training of the NN-based forward model, the posterior sampling of $\sigma$ plays a critical role in regularization. If $\sigma$ were held fixed during training, the model would be limited in its ability to adjust the sharpness of the likelihood surface and could fail to reach an optimal fit to the data. By treating $\sigma$ as a nuisance parameter and marginalizing over it within the BICePs framework, the model adaptively tunes the effective strength of the loss function, yielding an implicit form of regularization. This dynamic adjustment helps stabilize training and ensures that neural network parameters are optimized in a manner consistent with both the experimental data and the underlying physical ensemble.
[…]
Models have limited predictive reliability in regions lacking sufficient training data. A central advantage of our framework lies in its ability to incorporate physically relevant information through the choice of input features or explicit regularization. In our example using the toy model system, we employed the backbone dihedral angle $\phi$ as the sole input feature (projected to the unit circle to maintain periodicity). However, neural network inputs can encompass a broader range of features for complex observables such as chemical shifts, including quantum calculations, partial charges, descriptors of local atomic environments, interatomic distances, hydrogen bonding patterns, solvent accessibility, or torsional strain energies. These diverse features encode crucial structural and electronic context and can be precomputed from molecular simulations or quantum calculations before training.
We did not impose physical constraints or priors when training our model to predict $J$-coupling constants. Users might explicitly incorporate such physical priors or constraints through architectural decisions, regularization terms, or physically-informed input features, especially critical when dealing with sparse or unevenly distributed training data. For instance, in our experiments, neural networks employing Rectified Linear Unit (ReLU) activations struggled to accurately predict $J$-couplings at backbone dihedral angles greater than $+90^{\circ}$. In contrast, GELU activations provided smoother and more reliable extrapolation in these underrepresented regions. Thus, careful selection of activation functions and comprehensive input feature representations significantly enhances predictive accuracy across the full conformational landscape.
V Conclusion
@@
In the quest for accurate forward model predictions, specifically for $J$-coupling, researchers often navigate the vast literature seeking Karplus parameters that align with their specific systems, occasionally settling for less-than-ideal solutions. Our work demonstrates that BICePs can be used as a robust tool for determining forward model (FM) parameters by sampling over the full posterior distribution of their values and uncertainties.
We used a toy model of protein dihedral angles to demonstrate that posterior sampling and variational minimization of the BICePs score are valid and robust approaches for FM parameter refinement. Using structural ensembles and experimental data for ubiquitin, we applied BICePs with posterior sampling to optimize six different sets of Karplus coefficients using different types of $J$-coupling measurements, while effectively addressing both random and systematic errors.
From these results, one can see how BICePs can be applied more generally to optimize a wide variety of forward models, including those represented by neural networks. By treating the BICePs score as a physically grounded loss function, our framework enables the training of data-driven forward models that are consistent with ensemble-averaged experimental observables and incorporate principled uncertainty quantification. These advances not only contribute to the refinement of molecular simulations but also provide a scalable and extensible machine learning approach for physical sciences, with applications in structural dynamics, model validation, and the construction of predictive, interpretable models across a broad spectrum of data regimes.
@@
Data and software
The BICePs algorithm is freely available at github.com/vvoelz/biceps. All scripts, Jupyter notebook examples and analysis can be found here. Detailed documentation and tutorials can be found here: biceps.readthedocs.io. For any issues or questions, please submit the request on GitHub.
@@ Acknowledgements
The authors acknowledge that substantial content in this manuscript has also been included in the doctoral dissertation of RMR.Raddi (2024) RMR, TM and VAV are supported by National Institutes of Health grant R01GM123296. This research includes calculations carried out on HPC resources supported in part by the National Science Foundation through major research instrumentation grant number 1625061 and by the US Army Research Laboratory under contract number W911NF-16-2-0189.
Supplemental Information
Automatic Forward Model Parameterization with
Bayesian Inference of Conformational Populations
Robert M. Raddi, Tim Marshall and Vincent A. Voelz
@@
Appendix A: Second partial derivatives of the BICePs score with respect to forward model parameters.
The second partial derivatives of the BICePs score with respect to the FM parameters can be used for second-order optimization methods and uncertainty quantification. For simplicity, we refrain from showing the complicated second derivatives for the Good-Bad model in the general case. In the case when when $\varphi=1$, however, the second partial derivatives of the BICePs energy function with respect to parameters $\theta_{a}$ and $\theta_{b}$ are:
$\frac{\partial^{2}u}{\partial\theta_{m}\partial\theta_{n}}=N\left[\sum_{j=1}^{N_{d}}-\frac{\partial^{2}g_{j}(X,\theta)}{\partial\theta_{m}\partial\theta_{n}}\frac{(d_{j}-g_{j}(\mathbf{X},\theta))}{\sigma_{j}^{2}}\right.$ (S1)
$\left.+\frac{\partial g_{j}(X,\theta)}{\partial\theta_{m}}\cdot\frac{\partial g_{j}(X,\theta)}{\partial\theta_{n}}\frac{1}{{\sigma_{j}}^{2}}\right].$
When $m$ = $n$, the second partial derivative of the energy is just:
$\frac{\partial^{2}u}{\partial\theta_{m}^{2}}=N\left[\sum_{j=1}^{N_{d}}\left(-\frac{\partial^{2}g_{j}(X,\theta)}{\partial\theta_{m}^{2}}\frac{(d_{j}-g_{j}(\mathbf{X},\theta))}{\sigma_{j}^{2}}\right.\right.$ (S2)
$\left.\left.+\left(\frac{\partial g_{j}(X,\theta)}{\partial\theta_{m}}\right)^{2}\frac{1}{{\sigma_{j}}^{2}}\right)\right].$
The second partial derivatives of the BICePs score with respect to parameters $\theta_{m}$ and $\theta_{n}$ are:
@@
$\frac{\partial^{2}f(\theta)}{\partial\theta_{m}\partial\theta_{n}}=\bigg{\langle}\frac{\partial^{2}u}{\partial\theta_{m}\partial\theta_{n}}\bigg{\rangle}-\biggl{(}$ $\bigg{\langle}\frac{\partial u}{\partial\theta_{m}}\cdot\frac{\partial u}{\partial\theta_{n}}\bigg{\rangle}-\bigg{\langle}\frac{\partial u}{\partial\theta_{m}}\cdot\frac{\partial u}{\partial\theta_{n}}\bigg{\rangle}^{\prime}\biggr{)},$ (S3)
@@
where the notation $\langle\cdot\rangle^{\prime}$ denotes the ensemble average with respect to $\left(\frac{1}{Z}\exp(-u)\right)^{2}$. The first term on the right is the ensemble-average second derivative of the energy function $u$ with respect to parameters $\theta_{m}$ and $\theta_{n}$ given in equation S2 (when $\varphi=1$).
@@
When $m$ = $n$, the second partial derivative of the BICePs score reduces to the difference between the ensemble-averaged second derivative of the energy $u$ and the variance of its first partial derivative:
@@
$\begin{split}\frac{\partial^{2}f}{\partial\theta^{2}}&=\bigg{\langle}\frac{\partial^{2}u}{\partial\theta^{2}}\bigg{\rangle}-\left(\bigg{\langle}\left(\frac{\partial u}{\partial\theta}\right)^{2}\bigg{\rangle}-\bigg{\langle}\frac{\partial u}{\partial\theta}\bigg{\rangle}^{2}\right)\end{split}$ (S4)
@@
In practice, this calculation is performed using the MBAR free energy estimator for the BICePs score and it’s derivatives, by sampling at several intermediates $\xi=0\rightarrow 1$, which enables accurate estimates of all quantities.
Methods
Structural ensembles of human ubiquitin